Elton Company manufactures wheel rims. The controller expects the following The company expects to produce 500 units of eachRequirement 3. Compute the estimated ABC indirect manufacturing cost per unit of each model for 2018. Carry each cost to theActivity Materials handling Machine setup Insertion of parts Finishing Allocation Base Number of parts Number of setups NumbeStandard Deluxe 8.0 9.0 12.0 12.0 Parts per rim Setups per 500 rims Finishing hours per rim Total direct labor hours per rimacturing cos 1 - Actual overhead costs Actual qty of the allocation base used Predetermined overhead allocation rate Total es

expected data starts with parts per rim standard | deluxe
Answer #1
1 Total estimated overhead cost=Predetermined overhead allocation rate*Total allocation base
Materials handling:
Predetermined overhead allocation rate=$6 per part
Total allocation base=Total number of parts:
Standard Deluxe Total
Expected production a 500 500
Parts per rim b 8 9
Total number of parts c=a*b 4000 4500 8500
Total estimated overhead cost=6*8500=$ 51000
Machine setup:
Predetermined overhead allocation rate=$700 per setup
Total allocation base=Total number of setups
Standard Deluxe Total
Expected production a 500 500
Setups per 500 rims b 12 12
Total number of setups c=(a/500)*b 12 12 24
Total estimated overhead cost=700*24=$ 16800
Insertion of parts:
Predetermined overhead allocation rate=$23 per part
Total allocation base=Total number of parts=8500
Total estimated overhead cost=23*8500=$ 195500
Finishing:
Predetermined overhead allocation rate=$70 per hour
Total allocation base=Total direct labor hours:
Standard Deluxe Total
Expected production a 500 500
Finishing hours per rim b 6 6.5
Total number of finishing hours c=a*b 3000 3250 6250
Total estimated overhead cost=70*6250=$ 437500
Total estimated indirect manufacturing cost:
$
Materials handling 51000
Machine setup 16800
Insertion of parts 195500
Finishing 437500
Total 700800
2 Allocation rate=Total estimated indirect manufacturing cost/Total direct labor hours
Indirect manufacturing cost per rim:
Standard Deluxe Total
Expected production a 500 500
Total DLHs per rim b 7 8
Total number of DLHs c=a*b 3500 4000 7500
Allocation rate=700800/7500=$ 93.44 per DLH

a 93.44 b 4000 Indirect manufacturing cost per rim: Allocation rate Total number of DLHs Indirect manufacturing cost allocate